# Secret Sharing in the Rank Metric

### v1

@@ I Introduction
Linear rank-metric codes were originally introduced by Delsarte [8] as sets of matrices over a finite field $\mathbb{F}_{q}$, where the distance between two matrices is defined as the rank of their difference. These codes were later rediscovered by Roth [9] and Gabidulin [10]. Their properties have been widely studied in the context of random linear network coding [11, 12], including coset coding to secure the network from wiretappers [13, 14]. In addition, rank-metric codes have been proposed for post-quantum encryption and digital signature schemes. Notably, ROLLO and RQC are examples of rank-metric-based cryptosystem submissions to the NIST Post-Quantum Cryptography Standardization process [15, 16]. However, ROLLO and RQC did not advance in the PQC standardization process, as their security analyses required further scrutiny. Despite this, NIST recognizes that rank-based cryptography remains competitive and recommends continued research in this area [17]. As a result, rank-metric cryptanalysis remains an active area of research; see [18, 19, 20, 21]. Lastly, $q$-polymatroids, which generalize (poly)matroids, are well known to arise from rank-metric codes [22, 23, 24]. Consequently, $q$-polymatroids serve as a valuable tool to gain insights into rank-metric codes.
@@ II-A Access Structures and Secret Sharing
Next, let us briefly describe how Reed–Solomon codes yield secret sharing schemes. Let $0\leq k\leq n\leq q$ be integers and denote the vector space of polynomials over a finite field $\mathbb{F}_{q}$ with degree at most $k-1$ by $\mathbb{F}_{q}[X]_{<k}$. A Reed–Solomon code can be defined as
@@
Let $P$ be a finite set. Let $G,A\subseteq 2^{P}$ such that $G\cap A=\varnothing$, $G$ is monotone, and $A$ is an anti-monotone. Then $(G,A)$ is an access structure on $P$.
@@ III-A Access Structures on Subspaces
We prove the first equivalence as the second then follows by duality. Let $\pi\colon E\rightarrow E/Z$ denote the canonical quotient map. It is possible to choose bilinear forms such that $Z\oplus Z^{\perp}=E$ and there exists a monotone isomorphism $\phi\colon E/Z\rightarrow Z^{\perp}$ with a monotone inverse satisfying $\pi^{-1}(V)^{\perp}=\phi(V)^{\perp}\cap Z^{\perp}$ for all $V\in E/Z$, see [39, Theorem 5.3]. In particular,
$(\Gamma/Z)^{*}$ =$=\{V\in\mathcal{L}(E/Z)\,|\,\pi^{-1}(V^{\perp E/Z})\in\Gamma\},$
$\Gamma^{*}|_{Z^{\perp}}$ =$=\{W\in\mathcal{L}(Z^{\perp})\,|\,W^{\perp}\in\Gamma\},$
@@ III-B $q$-Polymatroid Ports
$\Gamma$ :-$\coloneq\{V\in\mathcal{L}(P)\,|\,\rho_{\mathcal{C}}(P_{0}\,|\,V)=0\},:$
$\mathcal{A}$$\coloneqq\{W\in\mathcal{L}(P)\,|\,\rho_{\mathcal{C}}(P_{0}\,|\,W)=\rho_{\mathcal{C}}(P_{0})\},$
@@
[…]
@@
$\mathbf{S}_{P_{0},P}(\mathcal{M})^{*}\simeq\mathbf{S}_{P^{\perp},P_{0}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1})}^{\perp}}(\mathcal{M}^{*}).$
@@
[…]
@@
$\mathbf{S}_{P_{0},P}(\mathcal{M}_{\mathcal{C}})$ is perfect as $10\leq\rho(P_{0}\,|\,V)\leq\rho(P_{0})=$ 1 for any $V\in\mathcal{L}(P)$. The secondary statement follows easily by definitions of restrictions and bases of $q$-matroids.
@@
This follows from the induced rank function of an MRD code being of the form $\rho_{\mathcal{C}}(V)=\min\{\dim V,\frac{\dim(\mathcal{C})}{m}\}$ for all $V\in\mathcal{L}(E)$ [22, Cor. 6.6]. ∎
@@ III-C The entropy function of representable $q$-polymatroids
Assume the hypothesis of Theorem 3. Let $V_{1},$ $\ldots,V_{\ell}\in\mathcal{L}(\mathbb{F}_{q}^{n})$. Then $H(\mathbf{Z}_{V_{1}},\ldots,\mathbf{Z}_{V_{\ell}})=H(\mathbf{Z}_{V_{1}+\cdots+V_{\ell}})$.
@@
Let $\bot=\bot\mathbb{F}_{q}^{n}$. Let $\sigma$ be the map $\sigma\colon\mathcal{C}\rightarrow\prod^{\ell}_{i=1}\mathcal{C}/\mathcal{C}(V_{i}^{\perp})$ defined by $\sigma(X)=(\pi_{V_{1}}(X),\ldots,\pi_{V_{\ell}}(X))$ for all $X\in\mathcal{C}$. This map is clearly $\mathbb{F}_{q}$-linear and satisfies $\ker\sigma=\cap^{\ell}_{i=1}\ker\pi_{V_{i}}=\ker\pi_{\sum_{i=1}^{\ell}V_{i}}$. The result then follows by the definition of joint entropy as $\mathrm{rk}\,\sigma=\mathrm{rk}\,\pi_{\sum^{\ell}_{i=1}V_{i}}$ and for any $Y=(Y_{1},\ldots,Y_{\ell})\in\mathrm{im}\,\sigma$ it holds that $\mathbb{P}(Y)=q^{\dim\ker\sigma-\dim\mathcal{C}}$. ∎
@@
Assume the hypothesis of Theorem 3. Let $V,W\in\mathcal{L}(\mathbb{F}_{q}^{n})$. Then $H(\mathbf{Z}_{W}\,|\,\mathbf{Z}_{V})=m\log(q)\rho_{\mathcal{C}}(W\,|\,V)$.

### v2

@@ I Introduction
Linear rank-metric codes were originally introduced by Delsarte [16] as sets of matrices over a finite field $\mathbb{F}_{q}$, where the distance between two matrices is defined as the rank of their difference. These codes were later rediscovered by Roth [40] and Gabidulin [22]. Their properties have been widely studied in the context of random linear network coding [29, 43], including coset coding to secure the network from wiretappers [44, Soljanin]. In addition, rank-metric codes have been proposed for post-quantum encryption and digital signature schemes. Notably, ROLLO and RQC are examples of rank-metric-based cryptosystem submissions to the NIST Post-Quantum Cryptography Standardization process [24, 23]. However, ROLLO and RQC did not advance in the PQC standardization process, as their security analyses required further scrutiny. Despite this, NIST recognizes that rank-based cryptography remains competitive and recommends continued research in this area [38]. As a result, rank-metric cryptanalysis remains an active area of research; see [4, 28, 9, 7]. Lastly, $q$-polymatroids, which generalize (poly)matroids, are well known to arise from rank-metric codes [27, 25, 42]. Consequently, $q$-polymatroids serve as a valuable tool to gain insights into rank-metric codes.
@@ II-A Access Structures and Secret Sharing
Next, let us briefly describe how Reed–Solomon codes yield secret sharing schemes. Let $0\leq k\leq n\leq q$ be integers and denote the vector space of polynomials over a finite field $\mathbb{F}_{q}$ with degree at most k1k-1 by $\mathbb{F}_{q}[X]_{<k}$. A Reed–Solomon code can be defined as
@@
Let $P$ be a finite set. Let $G,A\subseteq 2^{P}$ such that $\cap A=\varnothing, GG$ is monotone, and $A$ is an anti-monotone. Then $(G,A)$ is an access structure on $P$.
@@ III-A Access Structures on Subspaces
We prove the first equivalence as the second then follows by duality. Let $\pi\colon E\rightarrow E/Z$ denote the canonical quotient map. It is possible to choose bilinear forms such that $Z\oplus Z^{\perp}=E$ and there exists a monotone isomorphism $\phi\colon E/Z\rightarrow Z^{\perp}$ with a monotone inverse satisfying $\pi^{-1}(V)^{\perp}=\phi(V)^{\perp}\cap Z^{\perp}$ for all $V\in E/Z$, see [26, Theorem 5.3]. In particular,
$(\Gamma/Z)^{*}$ $=\{V\in\mathcal{L}(E/Z)\,|\,\pi^{-1}(V^{\perp E/Z})\in\Gamma\},$
$\Gamma^{*}|_{Z^{\perp}}$ $=\{W\in\mathcal{L}(Z^{\perp})\,|\,W^{\perp}\in\Gamma\},$
@@ III-B $q$-Polymatroid Ports
$\Gamma$ $\coloneq\{V\in\mathcal{L}(P)\,|\,\rho(P_{0}\,|\,V)=0\},$
$\mathcal{A}$ $\coloneqq\{W\in\mathcal{L}(P)\,|\,\rho(P_{0}\,|\,W)=\rho_{\mathcal{C}}(P_{0})\},$
@@
Let $\mathbf{S}=\mathbf{S}_{P_{0},P}(\mathcal{M})$ be a generalized $q$-polymatroid port and $\mathcal{P}(P)$ denote the $1$-dimensional spaces of $\mathcal{L}(P)$. Then $\sigma(\mathbf{S})=\max_{p\in\mathcal{P}(P)}\{\rho(p)\}/\rho(P_{0})$ is the information ratio of $\mathbf{S}$.
@@
$\mathbf{S}_{P_{0},P}(\mathcal{M})^{*}\simeq\mathbf{S}_{P^{\perp},P_{0}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1})}\hskip-4.2679pt^{\perp}}(\mathcal{M}^{*}).$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{ is independent for all }W<V}\}.$
@@
$\mathbf{S}_{P_{0},P}(\mathcal{M})$ is perfect as $0\leq\rho(P_{0}\,|\,V)\leq\rho(P_{0})=1$ for any $V\in\mathcal{L}(P)$. The secondary statement follows easily by definitions of restrictions and bases of $q$-matroids.
@@
This follows from the induced rank function of an MRD code being of the form $\rho_{\mathcal{C}}(V)=\min\{\dim V,\frac{\dim(\mathcal{C})}{m}\}$ for all $V\in\mathcal{L}(E)$ [27, Cor. 6.6]. ∎
@@ III-C The entropy function of representable $q$-polymatroids
Assume the hypothesis of Theorem 3. Let $V_{1},$ $\ldots,V_{\ell}\in\mathcal{L}(\mathbb{F}_{q}^{n})$. Then
$H(\mathbf{Z}_{V_{1}},\ldots,\mathbf{Z}_{V_{\ell}})=H(\mathbf{Z}_{V_{1}+\cdots+V_{\ell}}).$
@@
Let $\bot=\bot\mathbb{F}_{q}^{n}$. Let $\sigma$ be the map $\sigma\colon\mathcal{C}\rightarrow\prod^{\ell}_{i=1}\mathcal{C}/\mathcal{C}(V_{i}^{\perp})$ defined by
$\sigma(X)=(\pi_{V_{1}}(X),\ldots,\pi_{V_{\ell}}(X))$
for all $X\in\mathcal{C}$. This map is clearly $\mathbb{F}_{q}$-linear and satisfies $\ker\sigma=\cap^{\ell}_{i=1}\ker\pi_{V_{i}}=\ker\pi_{\sum_{i=1}^{\ell}V_{i}}$. The result then follows by the definition of joint entropy as $\mathrm{rk}\,\sigma=\mathrm{rk}\,\pi_{\sum^{\ell}_{i=1}V_{i}}$ and for any $Y=(Y_{1},\ldots,Y_{\ell})\in\mathrm{im}\,\sigma$ it holds that $\mathbb{P}(Y)=q^{\dim\ker\sigma-\dim\mathcal{C}}$. ∎
@@
Assume the hypothesis of Theorem 3. Let $V,W\in\mathcal{L}(\mathbb{F}_{q}^{n})$. Then
$H(\mathbf{Z}_{W}\,|\,\mathbf{Z}_{V})=m\log(q)\rho_{\mathcal{C}}(W\,|\,V).$